Grade 12Algebra

Give first the step you will use to separate the variable and then solve the equation:
4x = 25

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Now we have to divide both sides of the equation by 4,
Then we get,
= 4x/4 = 25/4
= x = 25/4

Thanks